# Helvipad: A Real-World Dataset for Omnidirectional Stereo Depth Estimation

### v1

@@ Abstract
Despite considerable progress in stereo depth estimation, omnidirectional imaging remains underexplored, mainly due to the lack of appropriate data. We introduce Helvipad, a real-world dataset for omnidirectional stereo depth estimation, consisting of 40K frames from video sequences across diverse environments, including crowded indoor and outdoor scenes with diverse lighting conditions. Collected using two 360° cameras in a top-bottom setup and a LiDAR sensor, the dataset includes accurate depth and disparity labels by projecting 3D point clouds onto equirectangular images. Additionally, we provide an augmented training set with a significantly increased label density by using depth completion. We benchmark leading stereo depth estimation models for both standard and omnidirectional images. The results show that while recent stereo methods perform decently, a significant challenge persists in accurately estimating depth in omnidirectional imaging. To address this, we introduce necessary adaptations to stereo models, achieving improved performance. Our dataset is available at https://vita-epfl.github.io/Helvipad/.
@@ 1 Introduction
Mobile robots and autonomous agents are increasingly used in real-world applications such as autonomous driving [16], healthcare [18], or agriculture [10], where accurate 3D representations of dynamic human environments are crucial for navigation and interaction. Depth estimation [34, 19, 32] aims to help build these representations by assessing the distance of surrounding objects from the ego agent.
@@
  • We propose and evaluate key adaptations for stereo matching models to better handle spherical image geometry, resulting in performances gains.
@@ 2 Related Work
Following early contributions from synthetic datasets like SceneFlow [7] and real-world benchmarks like KITTI [12, 23], large-scale driving datasets, such as Apolloscape [15], nuScenes [4], and Waymo [26] expanded the scope to more complex driving scenarios, but often lack pixel-wise depth labels or are limited to frontal views or multi-view configurations, as summarized in Tab. 1. In omnidirectional imaging, public datasets with 360° images have enabled deep learning methods to improve performance in tasks such as 2D and 3D object detection [3, 21, 38], semantic segmentation [3, 37], tracking [21], video summarization and monocular depth estimation [25]). OmniHouse and OmniThings [31] provide synthetic indoor scenes captured using a cross configuration of four fisheye cameras. Although this setup offers a wide field of view, it is more challenging and costly to implement than the simpler top-bottom setup of two 360° cameras in Helvipad. MP3D [5] and SF3D [3] offer 360° images but are limited to synthetic indoor scenes with single view. Wang et al. [28] built two stereo datasets with top-bottom setup based on these two synthetic databases. The JRDB dataset [21] spans a variety of tasks in real-world outdoor and indoor environments, but does not include pixel-wise depth annotations. In this context, Helvipad emerges as a comprehensive real-world stereo dataset of 360° images, covering indoor and outdoors scenes with varying lighting conditions and pixel-wise annotations.
@@
Deep learning-based methods have increasingly dominated stereo matching since the pioneering work of Zbontar and LeCun [35], who introduced the use of convolutional neural networks (CNNs) to describe image patches for stereo matching. Since then, many approaches have incorporated 3D convolutional kernels to construct cost volumes [17, 14], enhancing architectures with spatial pyramidal pooling, such as PSMNet [6]. Recently, many iterative optimization-based methods [19, 32, 36] proposed to recurrently refine the disparity field using local cost values. For instance, IGEV-Stereo [32] refines disparity maps iteratively using ConvGRUs and constructs a combined geometrical encoding volume that encodes both non-local geometry and fine-grained local matching details. Other methods use cross-view completion for self-supervised pre-training [29], a Markov Random Field with neural networks for potential functions and message passing [13], or an adapter to a pre-trained vision transformer for feature extraction [20].
@@
Unlike conventional stereo matching, research in omnidirectional stereo matching remains limited. Early models like OmniMVS [31] and SweepNet [30] introduced wide-baseline omnidirectional setups with fisheye images from multiple cameras, with SweepNet using spherical sweeping to generate dense cost volumes. Being the only approach to tackle omnidirectional stereo matching from top-bottom camera setup, 360SD-Net [28] is an end-to-end deep learning method that mitigates distortions in equirectangular images thanks to a polar angle coordinate input and a learnable cost volume. But its architecture with spatial pyramid pooling does not integrate more recent advances in stereo matching presented in the previous paragraph. In this paper, we enhance a recent leading stereo matching models with adaptations to better handle 360° images.
@@ 3.1 Data Acquisition
To assemble a dataset capturing human pedestrian activity, we developed a custom-built rig with which we collected data across distinct scenes, encompassing indoor (e.g., corridors, halls) and outdoor (e.g., squares, external parking, footpaths) settings in various lighting conditions on a university campus. The rig features two Ricoh Theta V 360° cameras in a top-bottom setup with a 19.1 cm baseline, capturing equirectangular images at 30 fps, paired with a Ouster OS1-64 LiDAR sensor with a vertical field of view of 42.4°, operating at 10 fps, and mounted 45.0 cm below the bottom camera. A central processor manages data capture and ensures device synchronization. In total, we collected 29 video sequences. More details on data collection are available in Appendix A.
@@ 3.2 Mapping LiDAR to 360° Images
The main challenge in processing our raw data is to ensure a robust and accurate projection of the LiDAR point clouds on the 360° camera images. The following is a summary of the algorithm’s approach, presented in  Algorithm 1). At the beginning of each recorded session, we select a set of key 3D points in the point cloud obtained by the LiDAR scans on a 19x19 chessboard, focusing on corners and edge midpoints to obtain their coordinates. The corresponding image pixel positions are selected in the equirectangular bottom image. We apply initial rotation and translation transformations $R_{\text{init}},T_{\text{init}}$ to make the camera lens the coordinate system’s center, ensuring alignment with the camera’s viewpoint: $p^{\text{cam}}_{i}=R_{\text{init}}p_{\text{LiDAR},i}+T_{\text{init}}$. The transformed points are then converted into spherical coordinates $(r,\theta,\phi)$ to match the structure of equirectangular images, and projected onto the equirectangular plane:
$x=\left(\frac{\phi+\pi}{2\pi}\right)\times\text{width},\quad y=\left(\frac{\theta}{\pi}\right)\times\text{height}.$ (1)
The core step of the mapping process is the optimization of the rotation and translation parameters, achieved by minimizing the error between the projected points and their corresponding location on the equirectangular images. Subsequently, these optimized rotation matrix and translation vector compose the best-fit transformation to align LiDAR point clouds on the bottom equirectangular images.
To assess alignment accuracy, we calculated the pixel error by measuring Euclidean distances between 200 manually selected LiDAR points and their image positions [24], resulting in an average error of 8.0 pixels, which represents 0.40% of the image diagonal, confirming our calibration quality. More details on the evaluation of the LiDAR point cloud mapping are available in Sec. B.3.
@@
7:     Project to equirectangular $(x^{\text{eq}}_{i},y^{\text{eq}}_{i})$ via $(\theta_{i},\phi_{i},W,H)$ using Eq. 1.
@@
In omnidirectional stereo imaging, spherical disparity is defined as the angular difference between 360° cameras (see Fig. 1). After mapping, we obtain disparity values from depth values to disparity via the following equation:
@@
(a) All sequences - Depth
(b) Indoor sequences - Depth
(c) Outdoor sequences - Depth
Figure 2: Histograms of depth values for the entire dataset without depth completion, indoor and outdoor sequences (day and night). Vertical dotted lines indicate the average depth for each setting.
@@ 3.3 Depth Completion from LiDAR Point Clouds
Temporal aggregation. To increase the number of valid points, we aggregate the point cloud of the current frame with those from the four preceding and four succeeding frames. Given this limited temporal window, the error introduced by the movement of the robot, or within the scene, is negligible, given the LiDAR’s high frame rate and the rig’s sub-pedestrian moving speed (see Sec. C.3).
@@
The relative term $\frac{r_{q}-r_{i}}{r_{q}}$ ensures that points across the entire depth range of the scene are treated equally. The choice of the uncertainty threshold is based on a predefined ratio of interpolated points (RIP) after filtering.
@@ 3.4 Benchmark and Data Statistics
After cropping unnecessary regions (sky, ground), the resulting dataset comprises 39,553 paired images with their corresponding depth and disparity maps, at a resolution of 1920 $\times$ 512 pixels. To create the benchmark, we split the captured sequences into a train-val set and a test set, ensuring the same proportion of outdoor, indoor and night outdoor sequences. The train-val set includes 20 sequences – 13 outdoor, 5 indoor and 2 night outdoor sequences – with a total of 29,407 frames. The test split consists of 6 sequences – 3 outdoor, 2 indoor and 1 night outdoor sequences – with a total of 10,146 frames.
Fig. 2 presents depth distributions of the entire dataset and by scene types. Depth values range from 0.5 to 225 meters, with averages of 8.1 meters overall, 5.4 meters for indoor scenes, and 9.2 meters for combined day and night outdoor scenes. Disparity distributions and distributions for augmented training data are also provided in Sec. A.2.
@@ 4 Adapting Stereo Matching for 360° Imaging
Method Type Disparity (°) Depth (m) Runtime (s)
PSMNet [6] stereo 0.33 0.54 0.20 2.79 6.17 0.29 0.66
360SD-Net [28] 360° stereo 0.21 0.42 0.18 2.14 5.12 0.15 0.63
IGEV-Stereo [32] stereo 0.22 0.41 0.17 1.85 4.44 0.15 $\bm{0.24}$
360-IGEV-Stereo 360° stereo $\bm{0.18}$ $\bm{0.39}$ $\bm{0.15}$ $\bm{1.77}$ $\bm{4.36}$ $\bm{0.14}$ 0.25
Table 2: Comparative results of depth estimation baselines on the Helvipad test set. All methods are trained on the augmented train set. 360-IGEV-Stereo achieves the strongest performance across most disparity and depth metrics. The best results are highlighted in bold.
@@ 5.1 Setup and Baselines
Evaluation metrics. We measure each method’s performance by comparing their disparity and depth estimates to the ground-truth labels using the Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), and Mean Absolute Relative Error (MARE). MAE indicates overall error magnitude, while RMSE captures the impact of outliers. In particular, depth errors are prone to increasing significantly due to the non-linear relationship between depth and disparity. MARE provides a scale-invariant measure of error, useful for assessing performance across varying depth ranges.
@@ 5.2 Comparative Results
Tab. 2 presents the performance of various stereo depth estimation methods evaluated on the Helvipad test set. While 360SD-Net improve every metrics over its rectilinear stereo counterpart PSMNet, it falls short of matching the depth accuracy of more recent models like IGEV-Stereo in metrics such as MAE and RMSE. This indicates that IGEV-Stereo has a notable advantage in minimizing larger depth errors compared to 360SD-Net. However, in terms of depth MARE and disparity metrics, 360SD-Net performs comparably to IGEV-Stereo, which suggests that 360SD-Net design for top-bottom 360° configurations helps to estimate accurately disparity.
With its omnidirectional adaptations, 360-IGEV-Stereo achieves the best performance across all disparity and depth metrics on the Helvipad test set. Compared to IGEV-Stereo, our adapted model reduces depth MAE from 1.85 to 1.77 meters and achieves a consistent improvement in disparity metrics, including a reduction of disparity MAE from 0.22° to 0.18°. Thus, highlighting benefits of adapting the stereo depth estimation baselines for omnidirectional data with a minimal overhead in computational cost.
@@ 5.3 Ablations
Depth completion as data augmentation. As shown in Fig. 4, training models with the depth-completed augmented set helps to achieve performance gains. For instance, 360SD-Net’s depth MARE dropped from 0.17 to 0.15 with augmentation, while PSMNet also shows a similar reduction. Notably, the performance gains are less pronounced for more advanced models like IGEV-Stereo and 360-IGEV-Stereo.
Figure 4: Depth MARE for different depth estimation methods trained with original or augmented depth labels.
Omnidirectional adaptations. During inference, circular padding is applied to account for the continuous 360° view, which helps to reduce the depth RMSE from 4.44 to 4.36, as shown in Tab. 3. This enhancement is visually noticeable in Fig. 6, where padding contributes to smoother depth transitions across the left and right image boundaries.
Further, the application of photometric augmentation helps models to generalize across the diverse lighting conditions in the Helvipad dataset.
Variant MAE $\downarrow$ RMSE $\downarrow$ MARE $\downarrow$
360-IGEV-Stereo $\bm{1.77}$ $\bm{4.36}$ $\bm{0.14}$
- w/o circular padding $1.78$ $4.39$ $\bm{0.14}$
- w/o photometric augmentation $1.80$ $4.38$ $\bm{0.14}$
Table 3: Ablation study on adaptations for 360-IGEV-Stereo method. The best results are highlighted in bold.
Figure 5: Cross-scene generalization analysis of each model when trained on different subsets (Indoor, Outdoor, Indoor+Outdoor, All) and evaluate under various testing conditions (Indoor, Outdoor, Night Outdoor). Performance are measured using depth MARE.

### v2

@@ Abstract
Despite progress in stereo depth estimation, omnidirectional imaging remains underexplored, mainly due to the lack of appropriate data. We introduce Helvipad, a real-world dataset for omnidirectional stereo depth estimation, featuring 40K video frames from video sequences across diverse environments, including crowded indoor and outdoor scenes with various lighting conditions. Collected using two 360° cameras in a top-bottom setup and a LiDAR sensor, the dataset includes accurate depth and disparity labels by projecting 3D point clouds onto equirectangular images. Additionally, we provide an augmented training set with an increased label density by using depth completion. We benchmark leading stereo depth estimation models for both standard and omnidirectional images. The results show that while recent stereo methods perform decently, a challenge persists in accurately estimating depth in omnidirectional imaging. To address this, we introduce necessary adaptations to stereo models, leading to improved performance.
@@ 1 Introduction
Mobile robots are increasingly used in real-world applications such as autonomous driving [18], healthcare [20], or agriculture [12], where accurate 3D representations of dynamic human environments are crucial for navigation and interaction. Depth estimation [38, 14, 41, 34] aims to help build these representations by estimating the distance of surrounding objects from the robot’s viewpoint.
@@
  • We propose and evaluate key adaptations for stereo matching models to better handle spherical image geometry, resulting in performance gains.
@@ 2 Related Work
Following early contributions from synthetic datasets like SceneFlow [9] and real-world benchmarks like KITTI [14, 26], large-scale driving datasets, such as Apolloscape [17], nuScenes [4], and Waymo [32] expanded to more complex scenarios, but often lack pixel-wise depth labels or are limited to frontal or multi-view configurations (see Tab. 1). In omnidirectional imaging, public datasets with 360° images have enabled deep learning methods to improve performance in tasks like 2D/3D object detection [3, 24, 45], semantic segmentation [3, 44], tracking [24], and monocular depth estimation [31]). OmniHouse and OmniThings [37] provide synthetic 360° indoor scenes captured using a cross-configuration of four fisheye cameras. Although this setup offers a wide field of view, it is more challenging and costly to implement than the simpler top-bottom setup of two 360° cameras in Helvipad. MP3D [5] and SF3D [3] offer 360° images but are limited to synthetic indoor scenes with single view. Wang et al. [34] built two stereo datasets with top-bottom setup based on these two synthetic databases. The JRDB dataset [24] spans a variety of tasks in real-world outdoor and indoor environments, but lacks pixel-wise depth annotations. In this context, Helvipad emerges as a comprehensive real-world stereo dataset of 360° images, covering indoor and outdoors scenes with varying lighting conditions and pixel-wise annotations.
@@
Deep learning has increasingly dominated stereo matching since Zbontar and LeCun [42] introduced the use of convolutional neural networks to describe image patches for stereo matching. Subsequent approaches have incorporated 3D convolutional kernels to regularize 4D cost volumes [19, 16] and enhanced architectures with spatial pyramidal pooling, like PSMNet [6]. Recently, iterative optimization-based methods [22, 38, 43] refine disparity fields recurrently using local cost values. For instance, IGEV-Stereo [38] refines disparity maps iteratively using ConvGRUs [8] and constructs a combined geometrical encoding volume that encodes both non-local geometry and fine-grained local matching details. Other methods use cross-view completion for self-supervised pre-training [35], a Markov Random Field with neural networks for potential functions and message passing [15], or an adapter to a pre-trained vision transformer for feature extraction [23].
@@
Unlike conventional stereo matching, research in omnidirectional stereo matching remains limited. Early models like OmniMVS [37] and SweepNet [36] introduced wide-baseline omnidirectional setups with fisheye images from multiple cameras, with SweepNet using spherical sweeping [27] to generate dense cost volumes. Recent works [28, 30] address spherical geometry distortions of panoramic images, while circular padding [21, 40] has been used to better maintain edge continuity in predictions. Unsupervised OmniMVS [7] leverages photometric consistency constraints to alleviate the need of omnidirectional stereo data, but may face challenges under low-light conditions and with non-Lambertian surfaces. Being the only approach to tackle omnidirectional stereo matching from top-bottom camera setup, 360SD-Net [34] is an end-to-end deep learning method that mitigates distortions in equirectangular images thanks to a polar angle coordinate input and a learnable cost volume. But its architecture with spatial pyramid pooling does not integrate more recent advances in stereo matching presented in the previous paragraph. In this paper, we enhance a recent leading stereo matching models with adaptations to better handle 360° images.
@@ 3.1 Data Acquisition
To assemble a dataset capturing human pedestrian activity, we built a custom rig and collected data across distinct scenes, encompassing indoor (e.g., corridors, halls) and outdoor (e.g., squares, external parking, footpaths) settings in various lighting conditions on a university campus. The rig features two Ricoh Theta V 360° cameras in a top-bottom setup with a 19.1 cm baseline, capturing equirectangular images at 30 fps, paired with a Ouster OS1-64 LiDAR sensor with a vertical field of view of 42.4°, operating at 10 fps, and mounted 45.0 cm below the bottom camera. A central processor manages data capture and ensures device synchronization. In total, we collected 29 video sequences. More details on data collection are available in Appendix A.
@@ 3.2 Mapping LiDAR to 360° Images
The main challenge in processing our raw data is to ensure a robust and accurate projection of the LiDAR point clouds on the 360° camera images. The following is a summary of the algorithm’s approach, presented in  Algorithm 1. At the beginning of each recorded session, we select a set of key 3D points in the point cloud obtained by the LiDAR scans on a 19x19 chessboard, focusing on corners and edge midpoints to obtain their coordinates. The corresponding image pixel positions are selected in the equirectangular bottom image. We apply initial rotation and translation transformations $R_{\text{init}},T_{\text{init}}$ to make the camera lens the coordinate system’s center, ensuring alignment with the camera’s viewpoint: $p^{\text{cam}}_{i}=R_{\text{init}}p_{\text{LiDAR},i}+T_{\text{init}}$. The transformed points are then converted into spherical coordinates $(r,\theta,\phi)$ to match the structure of equirectangular images, and projected onto the equirectangular plane. The core step of the mapping process is the optimization of the rotation and translation parameters, achieved by minimizing the error between the projected points and their corresponding location on the equirectangular images. Subsequently, these optimized rotation matrix and translation vector compose the best-fit transformation to align LiDAR point clouds on the bottom equirectangular images.
To assess alignment accuracy, we calculated the pixel error by measuring Euclidean distances between 200 randomly selected LiDAR points and their image positions [29], resulting in an average error of 1.7 pixels. When selecting the same number of points from visually challenging areas (e.g., object edges), the error remains relatively low at 8.0 pixels, further confirming calibration quality. More details on the evaluation of the LiDAR point cloud mapping are available in Sec. B.3. Occlusions were not explicitly handled, as the small LiDAR-camera baseline minimizes their impact.
@@
7:     Project to equirectangular: $(x^{\text{eq}}_{i},y^{\text{eq}}_{i})=\left(\frac{\phi_{i}+\pi}{2\pi}W,\frac{\theta_{i}}{\pi}H\right)$.
@@
In omnidirectional stereo imaging, spherical disparity is defined as the angular difference between 360° cameras-point rays (see Fig. 1). After mapping, we obtain disparity values from depth values to disparity via the following equation:
@@
(a) All sequences
(b) Indoor sequences
(c) Outdoor sequences
Figure 2: Histograms of depth values for the entire dataset without depth completion, indoor and outdoor (day and night) sequences. Vertical dotted lines indicate the average depth for each setting. Depth values range from 0.5m to 225m, with averages of 8.1m overall, 5.4m for indoor scenes, and 9.2m for combined day and night outdoor scenes.
@@ 3.3 Depth Completion from LiDAR Point Clouds
Temporal aggregation. To increase the number of valid points, we aggregate the the current frame’s point cloud with those from the four preceding and four succeeding frames. Given this limited temporal window, the error introduced by the movement of the robot, or within the scene, is negligible, given the LiDAR’s high frame rate and the rig’s sub-pedestrian moving speed (see Sec. C.3). Note that this approach is limited for fast-moving objects like cars.
@@
The relative term $(r_{q}-r_{i})/r_{q}$ ensures that points across the entire depth range of the scene are treated equally. The choice of the uncertainty threshold is based on a predefined Ratio of Interpolated Points (RIP) after filtering.
@@ 3.4 Benchmark and Data Statistics
After cropping unnecessary regions (sky, ground), the resulting dataset comprises 39,553 paired images with their corresponding depth and disparity maps, at a resolution of 1920 $\times$ 512 pixels. To create the benchmark, we split the captured sequences into a train-val set and a test set, ensuring the same proportion of outdoor, indoor and night outdoor sequences. Each test sequence has been manually reviewed to ensure no overlap of scenes or areas with the train set. The train-val set includes 20 sequences – 13 outdoor, 5 indoor and 2 night outdoor sequences – with a total of 29,407 frames. The test split consists of 6 sequences – 3 outdoor, 2 indoor and 1 night outdoor sequences – with a total of 10,146 frames. Fig. 2 presents depth distributions of the entire dataset and by scene types. Disparity distributions and distributions for augmented training data are also provided in Sec. A.2.
@@ 4 Adapting Stereo Matching for 360° Imaging
Method Stereo Setting Disparity (°) Depth (m) Runtime (s)
MAE $\downarrow$ RMSE $\downarrow$ MARE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$ MARE $\downarrow$ LRCE $\downarrow$
PSMNet [6] conventional 0.286 0.496 0.248 2.509 5.673 0.176 1.809 0.66
360SD-Net [34] omnidirectional 0.224 0.419 0.191 2.122 5.077 0.152 0.904 0.63
IGEV-Stereo [38] conventional 0.225 0.423 0.172 1.860 4.474 0.146 1.203 0.24
360-IGEV-Stereo omnidirectional 0.188 0.404 0.146 1.720 4.297 0.130 0.388 0.25
Table 2: Comparative results of depth estimation baselines on the Helvipad test set. All methods are trained on the augmented train set. 360-IGEV-Stereo achieves the strongest performance across all disparity and depth metrics. The best results are highlighted in bold.
@@ 5.1 Setup and Baselines
Evaluation metrics. We measure each method’s performance by comparing their disparity and depth estimates to sparse ground-truth labels using Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), and Mean Absolute Relative Error (MARE). MAE indicates overall error magnitude, while RMSE emphasizes the impact of outliers. Depth errors can increase significantly due to the non-linear relationship between depth and disparity, making MARE particularly useful as a scale-invariant metric for assessing performance across varying depth ranges. Dealing with 360° imaging, we also evaluate boundary consistency using the Left-Right Consistency Error (LRCE) [30].
@@ 5.2 Comparative Results
Tab. 2 presents the performance of various stereo depth estimation methods evaluated on the Helvipad test set. While 360SD-Net improve every metrics over its rectilinear stereo counterpart PSMNet, it falls short of matching the depth accuracy of more recent models like IGEV-Stereo in metrics such as MAE and RMSE. This indicates that IGEV-Stereo has a notable advantage in minimizing larger depth errors compared to 360SD-Net. However, in terms of depth MARE and disparity metrics, 360SD-Net performs comparably to IGEV-Stereo, which suggests that 360SD-Net design for top-bottom 360° configurations helps to accurately estimate disparity.
With its omnidirectional adaptations, 360-IGEV-Stereo achieves the best overall performance across all disparity and depth metrics on the Helvipad test set. Compared to IGEV-Stereo, our adapted model reduces depth MAE from 1.81m to 1.77m and achieves a consistent improvement in disparity metrics, including a reduction of disparity MAE from 0.22° to 0.18°. This highlights the benefits of adapting stereo depth estimation models for omnidirectional data with minimal computational overhead. As expected, 360-IGEV-Stereo also achieves the lowest LRCE, thanks to the use of circular padding during inference, and 360SD-Net, the other omnidirectional approach, ranks second.
@@ 5.3 Ablations
Depth completion as data augmentation. As shown in Fig. 4, training with the depth-completed augmented set improves performances. For instance, 360SD-Net’s depth MARE drops from 0.17 to 0.15 with augmentation, with PSMNet showing a similar reduction. The performance gains are less pronounced for more advanced models like IGEV-Stereo and 360-IGEV-Stereo.
Figure 4: Depth MARE comparison across different depth estimation methods trained on original vs. augmented depth labels.
Omnidirectional adaptations. During inference, circular padding is applied to account for the continuous 360° view, which helps to reduce the LRCE from 1.18m to 0.36m, and consequently also the depth RMSE, as shown in Tab. 3. This enhancement is visually noticeable in Fig. 6, where padding contributes to smoother depth transitions across the left and right image boundaries. Further, photometric data augmentation helps models to generalize across the diverse lighting conditions in the Helvipad dataset.
Variant MAE $\downarrow$ RMSE $\downarrow$ MARE $\downarrow$ LRCE $\downarrow$
360-IGEV-Stereo 1.720 4.297 0.130 0.388
- w/o circular padding 1.726 4.314 0.130 $1.153$
- w/o photometric augmentation $1.845$ $4.466$ 0.135 0.347
Table 3: Ablation study on adaptations for 360-IGEV-Stereo method. Results are reported using depth (m) metrics.
Figure 5: Cross-scene generalization analysis of each model when trained on different subsets (Indoor, Outdoor, Indoor+Outdoor, All) and evaluated under various testing conditions (indoor, outdoor, night outdoor). We use the depth MARE for comparison.
Figure 6: Visualization of disparity map predictions using IGEV-Stereo and 360-IGEV-Stereo compared to ground-truth sparse and augmented labels. The top row depicts images from the bottom camera, followed by the original and the augmented disparity maps from depth completion. The final two rows presents the predicted disparity maps from IGEV-Stereo and 360-IGEV-Stereo.